40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 94 of 348  Not logged in ELOG logo
ID Date Author Type Category Subject
  12808   Tue Feb 7 16:23:49 2017 SteveUpdateGeneralpower interruption tomorrow

                                                                                                                                   received this note: at 4:11pm Tuesday, Feb 7, 2017

**PLEASE POST**

 

Building:         Campus

    

Date:             Wednesday, February 8, 2017

          

Time:             7:30 AM – 8:30 AM  

 

Contact:          Rick Rodriguez x-2576

           

Pasadena Water and Power (PWP) will be performing a switching operation of the

Caltech Electrical Distribution System that is expected to be transparent to Caltech,

but could result in a minor power anomaly that might affect very sensitive equipment.

 

IMPACT: Negligible impact......?

There may be temporary  power interruption tomorrow!

PS:we did not see any effect   

  12807   Tue Feb 7 12:01:10 2017 LydiaUpdateIMC29.5 MHz stabilizer box replacement

I tested the amplifier with the Agilent network analyzer and measured 19.5 dB of gain between 29 and 30 mHz. The phase only changed by 1 degree over this same 1 MHz span. Since everything seems to be in order I'll hook it up this afternoon, unless there are any objections

Attachment 1: RF_amp.pdf
RF_amp.pdf
  12806   Tue Feb 7 10:18:58 2017 gautamUpdateIMCMC REFL weirdness

A few minutes back, I glanced up at the control room StripTool and noticed that the MCREFL PD DC level had gone up from ~0 to ~0.7, even though the PSL shutter was closed. This seemed bizzare to me. Strangely, simply cycling the shutter returned the value to the expected value of 0. I wonder if this is just a CDS problem to do with c1iool0 or c1psl? (both seem to be responding to telnet though...)

Since things look to be back to normal, I am going to start with my characterization of the various TFs in the IMC FSS loop...

  12805   Mon Feb 6 18:20:08 2017 KojiUpdateIMCMCL Feedback - simulink model updated

It is more accurate to model the physical frequency noises at various places.

cf. See also 40m ALS paper or Shigeo Nagano PDH thesis on https://wiki-40m.ligo.caltech.edu/40m_Library

- The output 4 should be "Laser frequency"

- Seismic path should be excluded from the summing node

- The output after the PMC: "Laser frequency after the PMC"

- "Laser frequency after the PMC" is compared (diffed) with the output 1 "mirror motion in Hz"

- The comparator output goes to the cav pole, the PD, and the PDH gain: This is the output named "PDH Error"

- Tap a new path from "Laser frequency after the PMC" and multiply with the cav pole (C_IMC)
- Tap a new path from "Mirror motion" and multiply with the cavity high pass  (s C_IMC/omega)
- Add these two: This is the output named "Frequency noise transmitted by IMC"

  12804   Mon Feb 6 17:03:41 2017 gautamUpdateIMCMCL Feedback - simulink model updated

I've edited Rana's Simulink model to reflect the current IMC servo topology (to the best of my understanding). I've tried to use Transfer Function blocks wherever possible so that we can just put in the appropriate zpk model in the script that will linearize the whole loop. I've also omitted the FSS SLOW loop for now.

I've been looking through some old elogs and it looks like there have been several modifications to both the MC servo board (D040180) and the TT FSS Box (D040105). I think it is easiest just to measure these TFs since the IMC is still down, so I will set about doing that today. There is also a Pomona Box between the broadband EOM and the output of the TT FSS box, which is meant to sum in the modulation for PMC locking, about which I have not yet found anything on the elog.

So the next steps are:

  1. Measure/estimate all the unknown TFs and gains in this schematic
  2. Linearize the model, get the OLG, see if the model matches previously measured OLGs (with the MCL part disabled)
  3. Once the model is verified to be correct, look at couplings of various noise sources in the MCL part of the loop, and come up with a suitable controller.

If anyone sees something wrong with this topology, please let me know so that I can make the required changes.

Attachment 1: mc40_v1.pdf
mc40_v1.pdf
  12803   Mon Feb 6 15:18:08 2017 gautamUpdateCDSslow machine bootfest

Had to reboot c1psl, c1susaux, c1auxex, c1auxey and c1iscaux today. PMC has been relocked. ITMX didn't get stuck. According to this thread, there have been two instances in the last 10 days in which c1psl and c1susaux have failed. Since we seem to be doing this often lately, I've made a little script that uses the netcat utility to check which slow machines respond to telnet, it is located at /opt/rtcds/caltech/c1/scripts/cds/testSlowMachines.bash.

The script can be executed by ./testSlowMachines.bash.

  12802   Mon Feb 6 10:05:28 2017 SteveUpdateSUSclamped cables

The bottom 5 cable connections from Sat-Amp to Whittering Filter at 1X5 were clamped today.

Attachment 1: clamped.jpg
clamped.jpg
  12801   Sun Feb 5 21:56:50 2017 LydiaUpdateIMC29.5 MHz stabilizer box replacement

Since the "stablizer box" doesn't really need to stabilize, it just needs to amplify, I decided to replace it with an off the shelf amplifier we already had, ZHL-2. I worked on getting it set up today, but didn't connect anything so that people have a chance to give some feedback. 

  • The gain we expect is 18 dB, and the maximum output with 1dB of compression is 29 dBm. To avoid compression, I'm aiming for ~26 dBm output, so ~8 dBm input. We measured the output of the source to be 12.8 dBm before, so I attached a 5dB attenuator to the input side of the amplifier. 
  • Across the 24V power input and the ground pin, I soldered a 100 uF, 50V electrolytic capacitor and a .27 uF, 50V metal film capacitor. Note that unlike the other similar amplifiers we have, the ground and +24 pins are separated (see image on datasheet). I wasn't sure if that changed what to do so I just found comparable caps to the ones that were there on another model. 
  • I twisted and soldered wires to the +24 and ground, making sure they were long enough to reach the clips where the power from the Sorensens gets split up. I placed the amplifer in the rack on top of the RF distribution box and ziptied the power cable in place. 
  • I connected a splitter to the output of the amplifier. Should I use a 10dB coupler instead, to maximize the power to the EOM?

So, I think the remaining thing to do is to connect the splitter to ASC out and to the line to the EOM, the +24V supply to the amplifier, and the 29.5 MHz input to the attenuator. I wanted to wait on this to get confiration that the setup is OK. Eventually we can put all of this in a box. 

Also, I noticed that in the clear cabinet with the Sorensens next to this rack, the +24 V unit is not supplying any voltage and has a red light that says "OVP." 

  12800   Sat Feb 4 12:50:01 2017 jamieSummaryCDS/cvs/cds/caltech/chans back on svn1.6
Quote:

No, not confused on that point. We just will not be testing OS versions at the 40m or running multiple OS's on our workstations. As I've said before, we will only move to so-called 'reference' systems once they've been in use for a long time.

Ubuntu16 is not to my knowledge used for any CDS system anywhere.  I'm not sure how you expect to have better support for that.  There are no pre-compiled packages of any kind available for Ubuntu16.  Good luck, you big smelly doofuses. Nyah, nyah, nyah.

  12799   Sat Feb 4 12:29:20 2017 jamieSummaryCDS/cvs/cds/caltech/chans back on svn1.6

No, not confused on that point. We just will not be testing OS versions at the 40m or running multiple OS's on our workstations. As I've said before, we will only move to so-called 'reference' systems once they've been in use for a long time.

Quote:
Quote:

True - its an issue. Koji and I are updating zita into Ubuntu16 LTS. If it looks like its OK with various tools we'll swap over the others into it. Until then I figure we're best off turning allegra back into Ubuntu12 to avoid a repeat of this kind of conflict. Once the workstations in the LLO control room are running smoothly on a new OS for a year, we can transfer into that. I don't think any of us wants to be the CDS beta tester for DV or DTT.

Just to be clear, since there seems to be some confusion, the SVN issue has nothing to do with Debian vs. Ubuntu.  SVN made non-backwards compatible changes to their working copy data format that breaks newer checkouts with older clients.  You will run into the exact same problem with newer Ubuntu versions.

I recommend the 40m start moving towards the reference operating systems (Debian 8 or SL7) as that's where CDS is moving.  By moving to newer Ubuntu versions you're moving away from CDS support, not towards it.

 

  12798   Sat Feb 4 12:20:39 2017 jamieSummaryCDS/cvs/cds/caltech/chans back on svn1.6
Quote:

True - its an issue. Koji and I are updating zita into Ubuntu16 LTS. If it looks like its OK with various tools we'll swap over the others into it. Until then I figure we're best off turning allegra back into Ubuntu12 to avoid a repeat of this kind of conflict. Once the workstations in the LLO control room are running smoothly on a new OS for a year, we can transfer into that. I don't think any of us wants to be the CDS beta tester for DV or DTT.

Just to be clear, since there seems to be some confusion, the SVN issue has nothing to do with Debian vs. Ubuntu.  SVN made non-backwards compatible changes to their working copy data format that breaks newer checkouts with older clients.  You will run into the exact same problem with newer Ubuntu versions.

I recommend the 40m start moving towards the reference operating systems (Debian 8 or SL7) as that's where CDS is moving.  By moving to newer Ubuntu versions you're moving away from CDS support, not towards it.

  12797   Sat Feb 4 12:00:59 2017 ranaSummaryCDS/cvs/cds/caltech/chans back on svn1.6

True - its an issue. Koji and I are updating zita into Ubuntu16 LTS. If it looks like its OK with various tools we'll swap over the others into it. Until then I figure we're best off turning allegra back into Ubuntu12 to avoid a repeat of this kind of conflict. Once the workstations in the LLO control room are running smoothly on a new OS for a year, we can transfer into that. I don't think any of us wants to be the CDS beta tester for DV or DTT.

  12796   Fri Feb 3 11:40:34 2017 ericqSummaryCDS/cvs/cds/caltech/chans back on svn1.6

I was able to bring back svn 1.6 formatting to /cvs/cds/caltech/chans by doing the following on nodus:

cd /cvs/cds/caltech
mkdir newchans
cd newchans
svn co https://nodus.ligo.caltech.edu:30889/svn/trunk/chans ./
rm -rf ../chans/.svn
mv ./.svn ../chans/

Note that I used the http address for the repository. The svn repository doesn't live at file:///cvs/cds/caltech/svn anymore; all of our checkouts (e.g. in the scripts directory) use http to get the one true repo location, regardless of where it lives on nodus' filesystem. (I suppose we could also use https://nodus.martian:30889/svn to stick to the local network, but I don't think we're that limited by the caltech network speed)

Presumably, at some point we will want to introduce a newer operating system into the 40m, as ubuntu 12.04 hits end-of-life in April 2017. Ubuntu 16.04 includes svn 1.8, so we'll also hit this issue if we choose that OS. 


Aside from the svn issues, this directory (/cvs/cds/caltech/chans) only contains pre-2010 channels. Filters and DAQ ini files currently live in /opt/rtcds/caltech/c1/chans, which is not under version control. It's also not clear to me why summary page configurations should be kept in this /cvs/cds place.

  12795   Fri Feb 3 11:40:09 2017 ranaUpdateIMCMCL Feedback - framing the problem

In working on automatic DARM loop design, we have this code:

https://git.ligo.org/rana-adhikari/ModernControls/tree/master/OptimalFeedback/GlobalCost

the things in there like mkCost*, etc. have examples of the cost functions that are used. It may be useful to look at those and then make a similar cost function calculation for the MCL/MCF loop.

  12794   Fri Feb 3 11:03:06 2017 jamieUpdateCDSmore testing fb1; DAQ DOWN DURING TEST

More testing of fb1 today.  DAQ DOWN UNTIL FURTHER NOTICE.

Testing Wednesday did not resolve anything, but Jonathan Hanks is helping.

  12793   Fri Feb 3 00:36:52 2017 gautamUpdateIMCMCL Feedback - framing the problem

Rana motivated me to take a step back and reframe the objectives and approach for this project, so I am collecting some thoughts here on my understanding of it. As I write this, some things still remain unclear to me, so I am leaving these as questions here for me to think about...

Objectives

  1. The PSL is locked to the IMC cavity - but at frequencies near 1 Hz, the laser frequency is forced to follow the IMC cavity length fluctuations, even though the free-running PSL frequency noise at those frequencies is lower. This excess is also imprinted on the arms when locked to the IR. We would like to improve the situation by feeding back a portion of the MC PDH error signal to the cavity length actuator to stabilize the MC cavity length at low frequencies. Moreover, we would like this loop to not imprint additional control noise in the arm control signals, which is a problem we have observed with the existing MCL loop. 
     
  2. The borader goal here is to use this project as a case study for designing the optimal loop and adaptive feedback. Can we come up with an algorithm, which takes
    • A model of our system (made with measured data where possible)
    • A list of our requirements (e.g. in this case, frequency noise requirements in various frequency bands, smooth crossovers between the various loops that enable locking the PSL to the IMC cavity and avoid injecting excess control noise into the plant)

and come up with the best loop that meets all our rquirements? What constitutes the "best" loop? How do we weight the relative importance of our various requirements? 


Proposed approach:

For the specific problem of making the MCL feedback loop better, the approach I have in mind right now is the following:

  1. Build a model of the 40m IMC loop. Ultimately the performance of the loop we implement will depend on the transfer function from various additive noise sources and disturbances in the feedback loop (e.g. electronics noise) to the output (i.e. laser frequency). Building an accurate model will allow us to quantify the performance of the proposed control loop, and hence, optimize it with some algorithm. I did some work on a simplistic, purely analytical model of the two MC loops (MCF and MCL), but Rana pointed out that it is better to have something more realistic for this purpose. I have inherited his Simulink models, which I will now adapt to reflect the 40m topology. 
  2. Come up with a list of requirements for the MCL controller. Some things that come to mind:
    • Reduce the arm control signal spectral amplitude below 20 Hz
    • Not increase the arm control signal spectral amplitude above 20 Hz
    • Crossover smoothly with the FSS slow temperature control loop and the MCF loop. 
    • What factor of suppression are we looking for? What is achievable? Once I build the model, it should shed some light on these..
    • Is the PMC a more stable frequency reference than the NPRO crystal at low frequencies? This measurement by Koji seems to suggest that it isn't (assuming the 1e4 product for the NPRO free-running frequency noise)..
  3. Once we have a model and a satisfactory list of requirements, design a control loop that meets these using traditional techniques, i.e. desired tracking error in the control band of 0.1-20 Hz (is this possible? The model will tell us...), gain and phase margin requirements etc. But this need not necessarily be the optimal controller that meets all of our requirements
  4. Optimize the controller - how? Can we define an objective function that, for example, rewards arm control signal suppression and penalizes injection of control noise, and just fminsearch in the [z,p,k] parameter space of the controller? Is there a smarter way to do this?
  5. Can this algorithm be adaptive, and optimize the controller to adapt to prevailing seismic conditions for example? Is this the same as saying we have a model that is accurate enough for us to predict the response of the plant to environmental disturbances? 

My immediate goal is to have the Simulink model updated.

Thoughts/comments on the above will be appreciated...

 
  12792   Thu Feb 2 18:32:51 2017 ranaSummaryPSLPMC alignment

Re-aligned the beam going into the PMC today around 5 PM. I noticed that its all in pitch and since I moved both of the mirrors by the same amount it is essentially a vertical translation.

I wonder if the PMC is just moving up and down due to thermal expansion in the mount? How else would we get a pure vertical translation? Need to remember next time if the beam goes up or down, and by how many knob turns, and see how it correlates to the lab temperature.

  12791   Thu Feb 2 18:28:29 2017 ranaSummaryCDSMinute Trend Koan

and the song remains the same...

the version of SVN on these workstations is ahead of the one on the other workstations so now we can't do 'svn up' on any of the Ubuntu12 machines. One allegra and optimus I get this error:

controls@allegra|GWsummaries> svn up
Updating '.':
svn: E180001: Unable to connect to a repository at URL 'file:///cvs/cds/caltech/svn/trunk/GWsummaries'
svn: E180001: Unable to open an ra_local session to URL
svn: E180001: Unable to open repository 'file:///cvs/cds/caltech/svn/trunk/GWsummaries'

Quote:
Quote:

Someone installed "Debian" on allegra. Why? Dataviewer doesn't work on there. Is there some advantage to making this thing have a different OS than the others? Any objections to going back to Ubuntu12?

My elog negligence punchcard is getting pretty full... It's pretty much for the same reason as using Debian for optimus; much of the workstation software is getting packaged for Debian, which could offload our need for setting things up in a custom 40m way. Hacking the debian-focused software.ligo.org repos into Ubuntu has caused me headaches in the past. Allegra wasn't being used often, so I figured it was a good test bed for trying things out.

The dataviewer issue was dataviewer's inability to pull the `fb` out of `fb:8088` in the NDSSERVER env variable. I made a quick fix for it in the dataviewer launching script, but there is probably a better way to do it.

I'm not sure if its possible to downgrade our chans repo back to the old one, but I highly recommend that no one do 'svn upgrade' in any of our repos until we remove all of the Debian installs in the 40m lab or hire a full-time sysadmin.

  12790   Thu Feb 2 17:43:20 2017 gautamUpdatePEMEM172 mic is hooked up in the PSL

I had noticed something wonky with the microphone, but neglected to elog it. I had tested it after installation by playing a sine wave from my laptop and looking at the signal on the PSL table, it worked fine. But you can see in the attached minute trend plot that the signal characteristics changed abruptly ~half a day after installation, and never quite recovered.\

 

Attachment 1: Mic_broken.png
Mic_broken.png
  12789   Thu Feb 2 17:34:25 2017 ranaUpdatePEMEM172 mic is hooked up in the PSL

I don't know if anyone looked at the time series (not trend) or spectrum of the Microphone after installation, but it looks bad and featureless to me. Is the Microphone broken?

This shows the spectrum from early this morning and again from tonight. You can see that it is bi-stable in its noise properties. This thing is busted; we're now removing it from the PSL so that it doesn' light it self on fire.

http://i.dailymail.co.uk/i/pix/2017/01/23/22/3C6BB70900000578-0-image-a-1_1485211840229.jpg

Attachment 1: MicFail.pdf
MicFail.pdf
  12788   Thu Feb 2 12:17:48 2017 SteveUpdateGeneral USB microscope

This AmScope microscope would have 3.5x-180x magnification, calibratable measurement function, 5MP picture and good working distance to work on printed circuit boards.

 

  12787   Thu Feb 2 11:25:45 2017 Max IsiUpdateSummary PagesCluster maintenance
FYI this issue has still not been solved, but the pages are working because I got the software running on an
alternative headnode (pcdev2). This may cause unexpected behavior (or not).

> LDAS has not recovered from maintenance causing the pages to remain unavailable until further notice.
> 
> > System-wide CIT LDAS cluster maintenance may cause disruptions to summary pages today. 
  12786   Wed Feb 1 23:13:30 2017 LydiaUpdateIMCRF AM stabilizer box Modification Plan

I made some of the changes. Gautam and I will finish tomorrow. 

While I was soldering the sharpest tip of the soldering iron (the one whose power supply shows the temperature) stopped working and I switched to a different one. Not sure how to fix this. 

Do we want to replace all of the removed ERA's with 50 Ohm resistors, or just the one along the spare output path? I shorted one of them with a piece of wire and left all the others open. 

I couldn't get one of the attenuators off (AT1, at beginning of ASC path). In trying I messed up the solder pad. Part of the connecting trace on the PCB board is exposed so we should be able to fix it. 

  12785   Wed Feb 1 20:49:34 2017 ranaUpdateIMCRF AM stabilizer box Modification Plan

I suggest:

  1. Disable the path which goes to the two spare outputs. Replace the ERA-5 with a 50 Ohm resistor to terminate that path. Make sure the ERA bias voltage is not shorting into something.
  2. Remove the ERA amps from the ASC path and remove the switch. Make it fixed gain such that we get +27 dBm out of the front.
  3. Put the ASC output into the 1U multi-splitter box and attenuate those outputs so that they supply ~0 dBm to the 2 WFS and the LSC Demod board.

I think this then allows us to have the low noise OCXO signals everywhere with enough oomph.

 

  12784   Wed Feb 1 16:45:56 2017 LydiaUpdateIMCRF AM stabilizer box Modification Plan

Here's what I'm planning to do to the RF AM stabilizer box. I'm going to take out several of the components along the path to the EOM (comments in green), including the dead ERA-4 and ERA-5 amplifiers, the variable attenuator which is controlled by a switch that can't be accessed outside the box, and the feedback path from the daughter board servo. I'm arranging things so that the output of the HELA-10 does not exceed the maximum output power. 

I wasn't quite as sure what to do about the path to the ASC box (comments in blue). I talked with Gautam and he said this gets split equally between several singals, one of which goes to the LO of the demod board which expects -10 dBm and currently gets -12 dBm (can go up to -8 by turning switch). So maybe we don't actually want the signal to be anywhere near +27 dBm at the output. The plans for the box are here, it looks like +27 in will end up with +10 at each output, which is way more than what's currently coming out. But maybe this needs to be increased to match the other path? 

Also we haven't measured the actual response of the variable attenuator U4 for various switch positions; it's the same model as the one I'm removing from the EOM path and that one had slightly different behavior for different switch positions than what the spec sheet says. Same goes for the HELA-10 units along this path: what is their actual gain? So perhaps these should be measured and then a single attenuator should be chosen to get the right output signal level. Alternatively it could just be left alone, if it is at an OK level right now. Advice on what to do here would be appreciated.  

I'll work on the EOM path tonight and wait for feedback on the rest of it. 

EDIT: Gautam pointed out that there's some insertion loss from the components I'll be removing that hasn't been accounted for. Also the plans have been updated to reflect that I'm replacing AT5 with a 1dB attenuator (from 6 dB). 

Attachment 1: RF_AM_stabilizer_modification.pdf
RF_AM_stabilizer_modification.pdf
  12783   Wed Feb 1 11:51:19 2017 KojiUpdateIMCRF AM stabilization box pulled out

For a comparison: OMC ELOG 238

  12782   Tue Jan 31 22:28:39 2017 LydiaUpdateIMCRF AM stabilization box pulled out

[rana, gautam, lydia]

Today we looked at the schematics for the RF AM stabilizer box and decided that there were an unnecessary amount of attenuators and amplifiers cancelling each other out and adding noise. At the end of the path are 2 HELA-10D amplifiers which we guessed based on the plots for the B version would have an acceptable amount of compression if the output of the second one is ~27dBm. This means the input to the first one should be a few dBm. This should be achieved with as simple a path as possible.

This begged the question, do we need the amplitude to be stabilized at all? Maybe it's good enough already when it comes into this box from the RF distribution box. So I tried to measure the AM noise of the 29.5 MHz signal that usually goes into the AM stabilizer:

  • I first measured the power to be 12.8 dBm with the AG4395.
  • I sent the signal through a splitter, then sent one side attenuated by 3 dB to the LO side of a level 7 mixer, and the other side attenuated by 10 dB to the RF side of the mixer.
  • The output of the mixer went through a lowpass filter at 1.9 MHz (with a 50Ω inline terminator). Initially I connected this directly to a DAQ channel (C1:ALS-FC_X_F_IN), but the ADC noise was stronger than the AM signal.
  • To fix this I used the SR560, AC coupled with a gain of 10^4. Attachment 1 is a spectrum of the noise measured with everything connected as described, and also for separate portions of the signal chain:
    • I measured the ADC noise by connecting a terminator to the cable going to DAQ.
    • I measured the mixer noise by putting a terminator on the RF input (and the end of the cable that was connected to it), while still driving LO.
    • I measured the SR560 noise by putting a terminator on the input.

It seems like I'm getting mostly noise from the SR560. Maybe it would be better to use an SR785 to take data instead of DAQ, and then skip the SR560? At low frequencies it seems like the AM noise measurement may be actually meaningful. In any case, if the actual AM noise from the crystal is lower than any of these other noise sources, it means we probably don't need to stabilize the amplitude with a servo, which means we can simplify the AM stabilizer board considerably to just amplify what it gets to 27 dBm.

Attachment 1: AM_noise.pdf
AM_noise.pdf
  12781   Tue Jan 31 22:15:02 2017 JohannesUpdateCDSvme crate backplane adapter boards

I made a crude sketch for how Lydia and I envision the connector situation on the back of the vme crates to be solved. Essentially the side panels of each crate extend about 2" (52 mm) beyond the edge of the DIN connectors. This is plenty of space for a simple PCB board. The connector of choice is D-Sub. We can split the 64 used pins into 2x 37 D-Sub OR (2x25 pin + 1x15pin). The former has fewer cables, but a few excess unused leads. A quick google search showed me that it is much cheaper to get twisted pair cables for 15 and 25 pin D-Subs. From what I remember, the used pins on the DIN connectors are concentrated on the low numbers end and the high numbers end, so might not need the 'middle' connector in many cases if we decide to break it up into three. I have to check this with Lydia though.

The D-Sub connectors would be panel mounted, for which we need a narrow panel piece with dsub cutouts. We can run horizontal struts across the vme crate from side panel to side panel. This way the force upon cable (dis)connection is mostly on the panel which is attached to the struts which are attached to the crate. This will also prevent gravitational sag or cable strain from pulling on the DIN connection, and we can use twisted pair cables with backshell, screws, and strain reliefs.

I was lookng into getting started with the PCB when Altium complained that the license is expired and to renew it. This is a relatively simple board layout so some free software out there is probably enough.

Attachment 1: vme_backplane_conn_sketch.jpg
vme_backplane_conn_sketch.jpg
  12780   Tue Jan 31 22:07:13 2017 gautamUpdateIMCRF AM stabilization box revamp

I've added the schematic of the RF AM stabilization board to the 40m PSL document tree, after having created a new DCC document for our 40m edits. Pictures of the board before and after modification will also be uploaded here...

  12779   Tue Jan 31 20:25:26 2017 ericqSummaryCDSMinute Trend Koan
Quote:

Someone installed "Debian" on allegra. Why? Dataviewer doesn't work on there. Is there some advantage to making this thing have a different OS than the others? Any objections to going back to Ubuntu12?

My elog negligence punchcard is getting pretty full... It's pretty much for the same reason as using Debian for optimus; much of the workstation software is getting packaged for Debian, which could offload our need for setting things up in a custom 40m way. Hacking the debian-focused software.ligo.org repos into Ubuntu has caused me headaches in the past. Allegra wasn't being used often, so I figured it was a good test bed for trying things out.

The dataviewer issue was dataviewer's inability to pull the `fb` out of `fb:8088` in the NDSSERVER env variable. I made a quick fix for it in the dataviewer launching script, but there is probably a better way to do it.

  12778   Tue Jan 31 18:51:07 2017 gautamUpdateSEISeismic Rainbow Strip - myths debunked

I've been suggesting that there may be something wonky with the Seismic Rainbow Striptool on the wall for the last couple of weeks. Here are a few things that were verified today.

  1. If you want to restore the StripTools in the control room, just run /opt/rtcds/caltech/c1/scripts/general/startStrip.sh. I have verified as of today that this works, and in future, any changes to channels/limits/colors of traces etc should be reflected in this script.
  2. Though some of the BLRMS bands have looked anomalous over the last few weeks, in particular the 0.3-1Hz band. The attached 120 day trend plot suggests that there hasn't been any dramatic change recently. In fact, looking on the summary pages, Rana noticed that today was an unusually low 0.3-1Hz activity day..
Attachment 1: Seis_BLRMS.png
Seis_BLRMS.png
  12777   Tue Jan 31 17:28:36 2017 ranaSummaryCDSMinute Trend Koan

Someone installed "Debian" on allegra. Why? Dataviewer doesn't work on there. Is there some advantage to making this thing have a different OS than the others? Any objections to going back to Ubuntu12?

  12776   Tue Jan 31 15:08:13 2017 ericqMetaphysicsCDSMinute Trend Koan

A novice was learning at the feet of Master Daqd. At the end of the lesson he looked through his notes and said, “Master, I have a few questions. May I ask them?”

Master Daqd nodded.

"Do we record minute trends of our data?"

"Yes, we record raw minute trends in /frames/trend/minute_raw"

"I see. Do we back up minute trends?"

"Yes, we back up all frames present in /frames/trend/minute"

"Wait, this means we are not recording our current trends! What is the reason for the existence of seperate minute and minute_raw trends?

“The knowledge you seek can be answered only by the gods.”

"Can we resume recording the minute trends?"

Master Daqd nodded, turned, and threw himself off the railing, falling to his death on the rocks below.

Upon seeing this, the novice was enlightened. He proceeded to investigate how to convert raw minute trends to minute trends so that historical records could be preserved, and precisely when Master Daqd started throwing himself off the mountain when asked to record minute trends.

  12775   Tue Jan 31 14:17:48 2017 gautamUpdateIMCRF AM stabilization box pulled out

> What is the probe situation? Ought to use a high impedance FET probe to measure this or else the scope would load the circuit.

We did indeed use the active probe, with the 100:1 attenuator in place. The values Lydia has quoted have 40dB added to account for this.

> What kind of HELA are the HELA amplifiers? Please a link to the data sheet if you can find it. I wonder what the gain and NF are at 30 MHz. I think the HELA-10D should be a good variant

The HELA is marked as HELA-10. It doesn't have the '+' suffix but according to the datasheet, it seems like it is just not RoHS compliant. It isn't indicated which of the varieties (A-D) is used either on the schematic or the IC, only B and D are 50ohms. For all of them, the typical gain is 11-12dB, and NF of 3.5dB.

  12774   Tue Jan 31 14:14:29 2017 ranaUpdateVACPressure watch script

I think this cron job is running on NODUS (our gateway) instead of our scripts machine:

*/1 * * * * /opt/rtcds/caltech/c1/scripts/Admin/n2Check.sh >> /opt/rtcds/caltech/c1/scripts/Admin/n2Check.log 2>&1

Quote:

Based on Jenne's chiara disk usage monitoring script, I made a script that checks the N2 pressure, which will send an email to myself, Jenne, Rana, Koji, and Steve, should the pressure fall below 60psi. I also updated the chiara disk checking script to work on the new Nodus setup. I tested the two, only emailing myself, and they appear to work as expected. 

The scripts are committed to the svn. Nodus' crontab now includes these two scripts, as well as the crontab backup script. (It occurs to me that the crontab backup script could be a little smarter, only backing it up if a change is made, but the archive is only a few MB, so it's probably not so important...)

moreover this script has a 90MB log file full of not finding its channel

I wish this script was in python instead of BASH and I wish it would run on megatron instead of nodus (why can't megatron send us email too?) and I wish that this log file would get wiped out once in awhile. Currently its been spitting out errors since at least a month ago:

Tue Jan 31 14:10:02 PST 2017 : N2 Pressure:

Channel connect timed out: 'C1:Vac-N2pres' not found.

(standard_in) 1: syntax error

  12773   Tue Jan 31 13:46:34 2017 ranaUpdateIMCRF AM stabilization box pulled out
  1. What is the probe situation? Ought to use a high impedance FET probe to measure this or else the scope would load the circuit.
  2. The ERA amplifiers are known to slowly die over ~10 year times scales. Search our ELOG for ERA-5. We'll have to replace some; ask Steve to order if we don't have many in the Plateau Tournant.
  3. What kind of HELA are the HELA amplifiers? Please a link to the data sheet if you can find it. I wonder what the gain and NF are at 30 MHz. I think the HELA-10D should be a good variant.
  12772   Tue Jan 31 01:07:20 2017 LydiaUpdateIMCRF AM stabilization box pulled out

[gautam, Lydia]

We looked at the RF AM stabilizer box to see if we could find out 1) Why the output power is so low, and 2) Why it can't be changed with the DC input "MOD CONT IN." Details to follow, attached is the annotated schematic from DCC document D000037

We are not returning the box tonight so the PSL shutter remains closed. 

Attachment 1: AM_stablilizer_annotation.pdf
AM_stablilizer_annotation.pdf
  12771   Mon Jan 30 19:07:48 2017 gautamUpdateIMCRF AM stabilization box pulled out

[johannes, gautam]

We pulled out the RF AM stabilization box from the 1X2 rack. PSL shutter was closed, marconi output, RF distribution box and RF AM stabilization box were turned off in that order. We had to remove the 4 rack nut screws on the RF distribution box because of the stiff cables which prevented the RF AM stabilization box extraction. I've left the marconi output and the RF distribution boxes off, and have terminated all open SMA connections with 50 ohm terminators just in case. Rack nuts for RF distribution box have been removed, it is currently sitting on a metal plate that is itself screwed onto the rack. I deemed this a stable enough ledge for the box to sit on in the short run, while we debug the RF AM stabilization box. We will work on the debugging and re-install the box as soon as we are done...

  12770   Mon Jan 30 18:41:41 2017 jamieUpdateCDSTEST ABORTED of new daqd code on fb1

I just aborted the fb1 test and reverted everything to the nominal configuration.  Everything looks to be operating nominally.  Front ends are mostly green except for c1rfm and c1asx which are currently not being acquired by the DAQ, and an unknown IPC error with c1daf.  Please let me know if any unusual problems are encountered.

The behavior of daqd on fb1 with the latest release (3.2.1) was not improved.  After turning on the full pipe it was back to crashing every 10 minutes or so when the full and second trend frames were being written out.  lame.  back to the drawing board...

  12769   Sat Jan 28 12:05:57 2017 jamieUpdateCDStest of new daqd code on fb1
Quote:

I'm not sure if this is related, but since today morning, I've noticed that the data concentrator errors have returned. Looking at daqd.log, there is a 1 second timing mismatch error that is being generated. Usually, manually running ntpdate on the front ends fixes this problem, but it did not work today.

If this problem started before ~4pm on Friday then it's probably unrelated, since I didn't start any of these tests until after that.  If unexplained problem persist then we can try shutting of the fb1 daqd and see if that helps.

  12768   Sat Jan 28 01:25:51 2017 gautamUpdateIMC29.5 MHz modulation depth

Some more details of our investigation:

  1. Here is a spectrum of the signal to the power combiner on the PSL table, measured on the output of the RF AM Stabilization box.

    Perhaps these sidebands were the ones I observed while looking at the input to the WFS demod board.
  2. The signal looked like a clean sinusoid when viewed on an oscilloscope with input impedance set to 50ohms. There were no sharp features or glitches in the time we observed, except when the 29.5 MHz MEDM slider was increased beyond 5, as noted by Lydia.
  3. We couldn't find a schematic for this RF AM Stabilization servo, so we are not sure what RF output power to the EOM we should expect. Schematic has since been found.
  4. I measured the power level at the input side (i.e. from the crystal) and found that it is ~12dBm, which seems reasonable (the front panel of the box housing the 29.5 MHz oscillator is labelled 13dBm). The schematic for the RF AM stabilization box says we should expect +10dBm at the input side, so all this points to a problem in the RF AM stabilization circuit...
  5. There is an attenuator dial on the front panel of the said RF AM stabilization servo that allows one to tune the power to the LO input of the WFS. Right now, it is set to approximately 7dB of attentuation, which corresponds to -12dBm at the WFS demod board input. I did a quick check to see if turning the dial changed the signal level at the LO input of the WFS board. The dial moves in clicks of 1dB, and the RF power at the LO input of the demod board increased/decreased by ~1dBm for each click the dial was rotated (I only explored the region 3dB-11dB of atttentuation). So it should be possible to increase the LO level to the WFS demod boards, is there any reason we shouldn't increase this to -8bBm (~0.25Vpp into 50ohms, which is around the level Koji verified the mixer to be working well at)?
  6. There were a couple of short ribbon cables which were just lying around on top of the cards in the eurocrate, Koji tells me that these were used as tester cables for checking the whitening filters and that they don't serve any purpose now. These have been removed.
  7. Added a button to IMC MEDM screen to allow easy access to the MEDM screen with slider to control the 29.5MHz modulation depth - though as mentioned in Lydia's elog, at the moment, this slider has no effect on the 29.5MHz power level to the EOM...
Attachment 1: IMC_mod.pdf
IMC_mod.pdf
  12767   Fri Jan 27 21:25:11 2017 LydiaUpdateIMC29.5 MHz modulation depth

[gautam, Lydia]

We set out to measure the 29.5 MHz power going to the EOM today but decided to start by looking at the output of the RF AM stabilizer box first. We wanted to measure the AM noise with a mixer, so we needed to know the power it was giving. We looked at the ouput that goes to the power combiner on the PSL table and found it was putting out only -2.0 dBm (~0.5 Vpp)! This was measured by taking a spectrum with the AG4395 and confirmed by looking on a scope.

To find out if this could be adjusted, we found an old MEDM screen (/opt/rtcds/caltech/c1/medm/c1lsc/master/C1LSC_RFADJUST.adl) and moved the 29.5 MHz EOM Mod Index Adjust slider while measuring the voltage coming in to the MOD CONTOL connection on the front of the AM stabilizer box. Moving the slider from 0 to 10 changes the input voltage linearly from -10 V to 10 V measured with a DMM at the cross-connects as we couldn't find an appropriate adapter for the LEMO cable. The 29.5 MHz modulation only appeared for slider values between 0 and 5, after which it abruptly shuts off. However, changing the slider value between 0 and 5 (Voltage from -10 to 0) does not change the amplitude of the output.

This seems like a problem; further investigation into the AM stabilizer box is neccessary. This DCC document outlines how to test the box, but we can't find a schematic. Since we don't have any mixers that can handle signals as small as -2 dBm, we gave up trying to measure the AM noise and will attempt to measure that and the reflection power from the EOM + resonant circuit once this problem has been diagnosed and fixed.

GV: After some digging, I found the schematic for the RF AM stabilization box (updated wiki and added it to the 40m document tree). According to it, there should be up to +22dBm of RF AM stabilized output to the EOM available, though we measured -2dBm yesterday, and could not vary this level by adjusting the EPICS voltage value. Neglecting losses in the cabling and the power combiner on the PSL, this translates to a paltry 0.178Vrms*0.6*8mard/Vrms ~ 0.85 mrad of modulation depth (gain at 29.5 MHz of the triple resonant circuit taken from this elog)... I think we need to pull this 1U chassis out and debug more thoroughly...

 

  12766   Fri Jan 27 21:21:35 2017 gautamUpdateCDSc1pem revamped

The coil and PD BLRMS are useful tools in identifying when glitches occur in the PD  readout, I thought it would be good to install them for ITMY, ETMX and SRM (since I plan to switch the MC3 satellite box, which we suspect to be problematic, with the SRM one). For this purpose, I had to install some IPC SHMEM blocks in C1SUS and recompile. 24 IPC channels were added to pipe the coil, PD and Oplev signals from C1SUS to C1PEM - the recompilation went smoothly, and it doesn't look like the model computation time has increased significantly or that the model is any closer to timing out.

However, I was unable to install the BLRMS blocks in C1PEM, as when I tried to compile the model with BLRMS for these extra 24 channels, I got a compilation error saying that I have exceeded the maximum allowed 499 testpoints per channel. Is there any workaround to this? It would be possible to create a custom BLRMS block that doesn't have all those testpoints, maybe this is the way to go? Especially if we want to install these channels for all our SOS optics, and also replace the current Seismic BLRMS with this scheme for consistency?

GV edit: I have implemented this scheme - after backing up the original BLRMS_2k part, I made a new one with no testpoints and only EPICS readouts. Doing so allowed me to recompile c1pem without any issues, the CPU time seems to have gone up by 3us from ~55us to 58us. So the BLRMS data record is only available at 16Hz, since there are no DQ channels in the BRLMS block - do we want these in any case? Let's see how this does over the weekend...

  12765   Fri Jan 27 20:52:36 2017 gautamUpdateCDStest of new daqd code on fb1

I'm not sure if this is related, but since today morning, I've noticed that the data concentrator errors have returned. Looking at daqd.log, there is a 1 second timing mismatch error that is being generated. Usually, manually running ntpdate on the front ends fixes this problem, but it did not work today.

Attachment 1: DCerrors.png
DCerrors.png
  12764   Fri Jan 27 19:40:03 2017 ranaMetaphysicselogword wrapping & large images

"Why does the word wrapping not work in our browsers with ELOG?" I sometimes wonder. Some of the elogs are fine, but often the 40m one has the text run off the page.

I found that this is due to people uploading HUGE images. If you need to do this, just use the shrink feature in the elog compose window so that we only have to see the thumbnail at first. Otherwise your 12 MP images will make it hard to read everyone else's entries.

  12763   Fri Jan 27 17:49:41 2017 jamieUpdateCDStest of new daqd code on fb1

Just FYI I'm running a test of updated daqd code on fb1. 

fb1 has it's own fiber to the daq network switch, so nothing had to be modified to do this test. This *should* not affect anything in the rest of the system, but as we all know these are famous last words....  If something is going haywire, and you can't get in touch with me and can't figure what else to do, you can just log on to fb1 and shut it down.  It's not writing any data to any of the network filesystems.

The daqd code under test is from the latest advLigoRTS 3.2.1 tag, which has daqd stability fixes that will hopefully address the problems we were seeing last time I tried this upgrade.  We'll see...

I'm going to let it run over the weekend, and will check in periodically.

  12762   Fri Jan 27 17:07:52 2017 LydiaUpdateCDSslow machine bootfest

Rebooted c1iscaux, c1auxex and c1auxey which were all not reponding to telnet. The watchdogs for the ETMs were turned off and then I keyed all 3 crates. All slow machines are reponding to telnet now. Both green lasers locked to the arms so I didn't do any burt restore.

  12761   Fri Jan 27 15:36:17 2017 KojiUpdateSUS wire standoffs update

Very nice! I got excited.

  • Don't you ask Calum and co to check the groove size with their microscopes? Just give the samples and the wire.
  • Do we want to make a simple "guitar" setup to measure the vibration Qs with Al piece, glass prism, ungrooved Sapphire, this grooved sapphire, grooved ruby, etc?
  12760   Fri Jan 27 14:50:04 2017 SteveUpdateSUS wire standoffs update

The 3 pieces of Sapphire v-groove test cuts are back. They look good. The suspension wire 0.0017" ( 43 micron ) is show on some of the pictures.

  12759   Fri Jan 27 00:14:02 2017 gautamSummaryIOOIMC WFS RF power levels

It was raised at the Wednesday meeting that I did not check the RF pickup levels while measuring the RF error signal levels into the Demod board. So I closed the PSL shutter, and re-did the measurement with the same measurement scheme. The detailed power levels (with no light incident on the WFS, so all RF pickup) is reported in the table below.

IMC WFS RF Pickup levels @ 29.5MHz
WFS Quadrant Pin (pW) Preflected
1 1 0.21 10.
2 1.41 148
3 0.71 7.1
4 0.16 3.6
2 1 0.16 10.5
2 1.48 166
3 0.81 5.1
4 0.56 0.33

These numbers can be subtracted from the corresponding columns in the previous elog to get a more accurate estimate of the true RF error signal levels. Note that the abnormal behaviour of Quadrant #2 on both WFS demod boards persists.

ELOG V3.1.3-